OpenStudy (anonymous):

What are the zeros of y=x(x-2)(x+5)?

OpenStudy (tylerd):

when x = 0 what happens?

OpenStudy (tylerd):

y=0(0-2)(0+5)=0

OpenStudy (tylerd):

theres still 2 more zero's u have to solve for, what do you think theyd be

OpenStudy (anonymous):

y=-10 when x=0?

OpenStudy (tylerd):

nope

OpenStudy (anonymous):

and do the same for y = 0?

OpenStudy (anonymous):

oh

OpenStudy (tylerd):

you have 0(0-2)(0+5) = 0(-2)(5)=0

OpenStudy (tylerd):

0 times -2 = 0

OpenStudy (anonymous):

oh right

OpenStudy (tylerd):

so you have to find all the x values when y=0

OpenStudy (tylerd):

x=0 is the first one

OpenStudy (anonymous):

x=2?

OpenStudy (anonymous):

x=-5?

OpenStudy (tylerd):

yep

OpenStudy (anonymous):

OK, got it now, thanks :D